x^3-27 over 5x^2+15x+45
show work

Answer by tutorcecilia(2144)

Apply the follow the rule for finding the difference of two cubes:
.
x^3 - a^3
= (x)^3 - (a)^3
= (x-a)(x^2 + ax + a^2)
.
x^3 - 27
= (x)^3 - (3)^3
=(x-3)(x^2 + 3x + 3^2)
= (x-3)(x^2 + 3x + 9)
Multiply the factors to check your answer.